OpenStudy (anonymous):

The ratio of an object's weight on Earth to its weight on Neptune is 5 : 7. How much would a person who weighs 150 pounds on Earth weigh on Neptune? A. 107 1/7 pounds B. 170 pounds C. 210 pounds D. 750 pounds

OpenStudy (anonymous):

5/7 = 150/x so solve for x x= 210

OpenStudy (austinl):

\(\dfrac{5}{7} = \dfrac{150}{x}\) Cross multiply and then divide. \(5x=150\times7\) \(x=\dfrac{150\times7}{5}\)

OpenStudy (anonymous):

Thank you! :3

OpenStudy (austinl):

You are welcome :)

OpenStudy (anonymous):

no problem!